# An Attention-Assisted AI Model for Real-Time Underwater Sound Speed Estimation Leveraging Remote Sensing Sea Surface Temperature Data

### v2

@@ IV-B Interpretability analysis
To investigate why the proposed SA-MDF-CNN model has better performance compared to CNN, we conducted an experiment and visually represented the attention weights of the model as shown in figure 11. Figure 11 (a) illustrates the first three order EOF feature vectors of the historical SSPs at various latitude and longitude coordinates. Figure 11 (b), (c), and (d) correspond to the attention weights when the training epochs are 10, 50, and 100, respectively. Through observing the evolution trend of attention focus with increasing training times, we found that the SA-MDF-CNN model increasingly focuses on the sound velocity distribution in shallow waters within 1000 meters. This trend aligns with the first three orders of EOF feature distribution shown in panel (a), although the EOF itself emphasizes the SSP distribution between 0-500 meters. The attention depth of the proposed model spans 0-1500 meters, which is attributed to the variation of SSP distribution under different latitude and longitude coordinates. Therefore, in order to enhance the model’s ability to focus on the global SSP distribution, we added a multi-head self-attention layer after input data fusion. This mechanism further confirms the effectiveness of the proposed model. Experiments based on the aforementioned public dataset demonstrate that the proposed SA-MDF-CNN model outperforms other algorithms in estimating sound velocity in both shallow and deep waters.
@@ IV-C Efficiency and stability
Although the accuracy performance in SSP estimation of SA-MDF-CNN is better than traditional methods, the disadvantage is that SA-MDF-CNN increases the number of parameters and computational overhead. Table IV shows the comparison of parameters and running time between the proposed model and CNN. It can be seen that the proposed model in this paper needs more parameters and training time than CNN, which is a necessary overhead for performance improvement.
@@ V Ocean experiments
Figure 14: Comparison of real-time SSP estimation results of different algorithms on ocean experimental data, where (a) shows the SSP curves, and (b) shows the error distributions.
In order to verify the practicality of the SA-MDF-CNN model, a comprehensive deep-ocean experiment was conducted in the South China Sea in April 2023. As illustrated in figure 13 (a), various instruments such as CTD and XCTD were employed at diverse marine locations, yielding a total of 14 distinct samples. The depth of sampling reached 3500 meters, with intervals set at every 100 meters. It is evident that the selected location for this experiment does not conform to the rigid 3x3 grid pattern, thus allowing for a more rigorous evaluation of the proposed model. The fusion data at the blue coordinates in figure 13 (b) were selected as the input of the model, and the real-time SSP under the red pentagram was used as the prediction output of the model.
Figure 15: Comparison of real-time estimation SSP results of different algorithms in shallow sea environment with different depths in ocean experiment, where (a) and (c) are SSP curves of 300 meters depth, (b) and (d) are error distributions of 500 meters depth.
@@
To evaluate the effectiveness of the model in estimating the distribution of sound velocity at full-ocean depth, the estimated SSP of SA-MDF-CNN is compared with other algorithms in figure 14. It is evident that the proposed model exhibits lower RMSE and more stable error disturbance compared to traditional algorithms and CNN. To further test the sound velocity estimation performance in shallow waters with significant changes in sound velocity, estimation of sound velocity at 300, 500 depth meters by SA-MDF-CNN are compared with other methods in figure 15. The results indicate that the proposed algorithm has smaller error disturbances and maintains significant accuracy advantages compared to other algorithms in shallow water environments. Figure 16 gives a more detailed comparison of the RMSE results of different algorithms, clearly showing that the proposed model has significant advantages at different depths.
@@ VI Conclusion
To construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, a SA-MDF-CNN model driven by multimodal data fusion is proposed in this paper, which fuses historical SSP features, real-time remote sensing SST, latitude and longitude coordinates. The fusion data is constructed by gridding, and the real-time SSP is retrieved based on the real-time remote sensing SST. The effectiveness of our proposed model was tested in the case of non-standardized grid through ocean experiments. The experimental results show that our proposed model is not only suitable for real-time SSP estimation under non-strictly standardized grid conditions, but also suitable for shallow sea conditions with different depths.
@@ Acknowledgments
The authors acknowledge the historical SSP data support from the China Argo Real-time Data Center, (https://www.argo.org.cn/, latest access:December 20, 2024). The remote sensing SST data can be acquired at the National Oceanic and Atmospheric Administration (https://www.commerce.gov/, latest access: December 25, 2024).

### v3

@@ IV-B Interpretability analysis
To investigate why the proposed SA-MDF-CNN model has better performance compared to CNN, we conducted an experiment and visually represented the attention weights of the model as shown in figure 11. Figure 11 (a) illustrates the primary three-order EOF eigenvectors corresponding to historical SSPs at multiple geospatial coordinates. Figure 11 (b), (c), and (d) correspond to the attention weights when the training epochs are 10, 50, and 100, respectively. Through observing the evolution trend of attention focus with increasing training times, we found that the SA-MDF-CNN model increasingly focuses on the acoustic velocity profile in shallow marine environments within 1000 meters. This trend aligns with the first three orders of EOF feature distribution shown in figure 11 (a), although the EOF itself emphasizes the SSP distribution between 0-500 meters. The attention depth of the proposed model spans 0-1500 meters, which is attributed to the variation of SSP distribution under different latitude and longitude coordinates. To improve the capability of capturing global SSP distribution patterns, a multi-head self-attention mechanism was incorporated following the fusion of input data. This architectural enhancement demonstrates the robustness and validity of the introduced framework. Experiments based on the aforementioned public dataset demonstrate that the proposed SA-MDF-CNN model outperforms other algorithms in estimating sound velocity in both shallow and deep waters.
@@ IV-C Efficiency and stability
Although the accuracy performance in SSP estimation of SA-MDF-CNN is better than traditional methods, the disadvantage is that SA-MDF-CNN increases the number of parameters and computational overhead. Table IV presents the computational complexity analysis and temporal efficiency comparison between SA-MDF-CNN and conventional CNN architectures. The experimental results indicate that the introduced methodology requires increased parameter quantities and extended training durations compared to CNN, representing an essential computational investment for achieving enhanced predictive capabilities.
@@ V Ocean experiments
In order to verify the practicality of the SA-MDF-CNN model,a deep-sea experiment was conducted in the South China Sea in April 2023. As illustrated in figure 13, various instruments such as CTD and XCTD were employed at diverse marine locations, yielding a total of 14 distinct samples. The depth of sampling reached 3500 meters, with intervals set at every 100 meters. It is evident that the selected location for this experiment does not conform to the rigid 3x3 grid pattern, thus allowing for a more rigorous evaluation of the proposed model. The input data for the model were extracted from the fusion dataset corresponding to the blue coordinates, while the real-time SSP associated with the red pentagram served as the model’s predictive output.
Figure 14: Evaluation of real-time SSP estimation outcomes utilizing diverse algorithms on oceanographic experimental data, where (a) shows the SSP curves, and (b) shows the error distributions.
Figure 15: Comparison of real-time estimation SSP results of different algorithms in shallow sea environment with different depths in ocean experiment, where (a) and (b) are SSP curves of 300 meters depth, (c) and (d) are error distributions of 500 meters depth.
To evaluate the effectiveness of the model in estimating the distribution of sound velocity at full-ocean depth, the estimated SSP of SA-MDF-CNN is compared with other algorithms in figure 14. It is evident that the proposed model exhibits lower RMSE and more stable error disturbance compared to traditional algorithms and CNN. To further test the sound velocity estimation performance in shallow waters with significant changes in sound velocity, estimation of sound velocity at 300, 500 depth meters by SA-MDF-CNN are compared with other methods in figure 15. The experimental outcomes demonstrate that the introduced methodology exhibits reduced error fluctuations and sustains notable precision benefits relative to alternative approaches in shallow marine settings. This comparative analysis highlights the robustness and superior performance of the proposed technique under such environmental conditions. Figure 16 gives a more detailed comparison of the RMSE results of different algorithms, clearly showing that the proposed model has significant advantages at different depths.
@@ VI Conclusion
To construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, an SA-MDF-CNN model driven by multimodal data fusion is proposed in this paper, which fuses historical SSP features, real-time remote sensing SST, latitude and longitude coordinates. The attention mechanisms were applied to enhance the model’s capacity to learn complex interactions within multimodal data. To evaluate the accuracy performance of the model, we conducted experiments using Argo data from the Pacific region and further validated it through non-standardized grid ocean experiments in the South China Sea in April 2023. The empirical findings demonstrate that the developed framework exhibits dual applicability, functioning effectively for real-time SSP reconstruction in non-uniform grid environments while maintaining effectiveness across diverse shallow-water and deep-water scenarios. At the same time, in order to enhance the interpretability of the model, we visualized the model’s parameters and found that the parameter weights were more concentrated in shallow water parts, which is consistent with the magnitude of the change in sound velocity. SA-MDF-CNN provides a fast and accurate way to estimate the distribution of regional sound velocity, which is suitable for areas with stable marine environments. However, it still has some limitations, for example, the model is insensitive to fluctuations in sound velocity distribution caused by internal ocean waves. Therefore, in future work, we will focus on studying the accurate estimation of sound velocity distribution under the influence of ocean internal waves without conducting on-site measurements of underwater data.
@@ Acknowledgments
The authors acknowledge the historical SSP data support from the China Argo Real-time Data Center, (https://www.argo.org.cn/, latest access:December 20, 2024). The remote sensing SST data can be acquired at the National Oceanic and Atmospheric Administration (https://www.commerce.gov/, latest access: December 25, 2024). There is no conflict of interest.